# From the light curves of Long Period Variables to their evolution along the AGB

### v1

@@ 3.5 Humpy curves and hump-less curves
In Paper I, we remarked that when a hump was observed on the ascending branch, there was a relation between its location and the parameters , and : namely the closer it was from minimal light, the larger were the values of these three parameters. Figure 8 illustrates the present situation. We selected curves displaying a clear hump on the ascending branch and classified them as type b, c or d according to the hump location: b for humps close to minimal light, d for humps close to maximal light and c in between. In the definitions of both what we call a strong hump and of which class the curve belongs to, there is of course some subjectivity which must be kept in mind. Moreover, what we define as humps are rather abrupt changes of the rate of increase of the luminosity: while c types are real humps, b and d types are rather flattening of the light minimum and maximum, respectively. Figure 8 reveals a number of remarkable features: all d types are from the Table 4 sample and, with the exception of V Cyg, all b types are from the Table 3 sample; instead, c types are present in all three samples and span a very large range of values; the general trends observed in Paper I are clearly confirmed and complemented with the dependence of the hump type on , and . Such humps may simply be the trivial effect of non-linear dynamics, similar humps are observed in RR Lyrae stars and in bump Cepheids with the Hertzsprung progression. In such a case, they may not be amenable to a simple interpretation, such as tracking a shock wave as suggested by L. S. Kudashkina & G. M. Rudnitskij (1994). Yet, their distribution in the parameter space suggests that they carry some additional information, which we however fail to decrypt.
@@ 4.1 Hot Bottom Burners
R Cen and R Nor have been identified as clear Hot Bottom Burning (HBB) candidates at the beginning of the TP-AGB by D. A. Garc´ıa-Hernández et al. (2013) and S. Uttenthaler et al. (2011, 2012) on the basis of detailed evaluations of the abundances of technetium, lithium, rubidium and zirconium in the spectra: they are Tc-poor and Li-rich and have low mass-loss rates at the level of 10-7 M⊙ yr-1, properties which they share with two other clear HBB candidates, SV Cas and RU Cyg, both semi-regular variables. They are expected to have initial masses of 4 to 6 solar masses, in contrast to more massive HBB candidates with very high mass-loss rates (10-5 M⊙ yr-1), known as OH/IR stars, at the end of the TP-AGB (D. A. Garc´ıa-Hernández et al., 2006, 2007, 2009). Their light curves are double-peaked (Figure 9) and that of R Cen has recently experienced a sudden change in period and amplitude. Moreover, they are spectacular outliers in colour vs period diagrams (S. Uttenthaler et al., 2019; P. Merchán Benítez et al., 2023; P. Whitelock et al., 2000). They have been commented upon in numerous earlier studies, both for R Cen (P. C. Keenan & J. Landi Dessy, 1966; G. Hawkins et al., 2001; W. S. G. Walker & J. Greaves, 2001; M. R. Templeton et al., 2005; J. R. Garcia & S. A. Dominguez, 2006; N. Vogt et al., 2016; B. Rea & J. C. Martin, 2024) and for R Nor (N. Vogt et al., 2016). Most authors take it as granted that the sudden changes observed in the light curve of R Cen are the result of a recent thermal pulse, as had been first proposed by P. R. Wood & D. M. Zarro (1981). Measurements of the radial velocity in the near infrared (T. Lebzelter et al., 2005) have given evidence for a clear anti-correlation with the light curve.
@@ 4.2 Mno outliers
The curve of W Dra is only marginally different from those of the Table 1 sample, in which it could be included: we did not find in the published literature sufficient reason to give it a special treatment.
@@ 5 Discussion and conclusion
The emphasis in the present article has been to try to reveal general relations between the evolution of stars along the AGB and the parameters of their light curves, at the price of accepting some loss of rigour. For example, by using mean parameters , and , we lost the information contained in the differences between the members of the relevant pairs. Also, we accepted to ignore the differences, often significant, between the uncertainties attached to the parameters that have been used. This must be kept in mind when considering the results obtained and possibly extending the present work to future observations or analyses. By using five parameters to characterize the main features of a light curve, we have been able to clarify the picture that had been sketched in Paper I. Figure 10 and Table 7 illustrate the main results, which we summarize below. While on the early part of the AGB, stars expand, cool down and tend to pulsate with increasing periods; when their period exceeds 100 days, they are classified as Mira variables and enter the sample considered in the present work. At that time, they are oxygen-rich with an M spectral type and have not yet experienced sufficient TDU events for their impact to be detectable. The sample of Mno curves listed in Table 1 corresponds to such stars. As the temperature of the star decreases and as its size increases, the light curve evolves from what was defined as first Mno family to what was defined as second Mno family in Paper I. Namely (black arrow in Figure 10) the period increases, the width parameter decreases, the phase at minimal light increases, weak humps may appear on the ascending branch close to minimal light, the pulsations become less regular, the colour indices increase and the ratio between the oscillation amplitudes in the infrared and in the visible increases.
@@
What the present work has made clear is that the transition from Mno to Myes, namely the occurrence of TDU events strong enough to have a detectable impact, occurs at very different stages along the TP-AGB. Indeed it is expected to never occur for stars of low initial masses (F. Herwig, 2005), and the curves of such stars stay in the Table 1 sample for the whole AGB life of the star. The presence in the Table 3 sample of stars of Myes, and even S spectral types covering a broad spectrum of values is compatible with the expectation that TDU events occur the earlier on the TP-AGB, the larger the initial mass of the star. The curves of the Table 3 sample (blue arrow in Figure 10) seem to keep evolving in the same direction after the transition to the Myes and S spectral types: the periods and amplitudes of the oscillations keep increasing and they become less regular. This is consistent with the expectation that thermal pulses modify only progressively the internal state of the star. The absence of carbon-rich curves in the parameter space of the curves close to the more evolved curves of the Table 3 sample suggests that some of these curves leave the domain of stable pulsation and become chaotic before the transition from oxygen-rich to carbon-rich.
@@
Another major result obtained by the present work is the existence of curves leaving the Table 1 sample very early and transiting to the Myes and S spectral types with the width parameter increasing rather than decreasing. These are the curves of the Table 4 sample and their location in the parameter space suggests that they will evolve to the Table 5 sample when the stars will become carbon-rich (red arrows in Figure 10). Here again, such a suggestion, albeit sensible and plausible, is far from being proven. The dispersion of the curve parameters about their mean is large for both the Table 4 and Table 5 samples, showing the complexity of the situation. In particular, the presence in the Table 5 sample of curves evolving with decreasing (possible successors of the Table 3 sample) and of curves evolving with increasing (possible successors of the Table 4 sample) contribute to such complexity. It is natural to speculate that the stars of the Table 4 sample have larger initial masses, and probably are closer to exhausting their convective layers than the stars of the Table 3 sample: the behaviour of the curves of HBB stars discussed in Subsection 4.1 gives support to such speculation as do the curves of stars that are known to be close to post-AGB, such as V Hya and VX Sgr.
@@
A major difficulty that has been recurrently met during the study is the lack of a clear picture of the physics governing the transition from a stable to an unstable regime of pulsation, namely of the nature of the stability strip in the Hertzsprung-Russell diagram. The distinction between Mira, semi-regular and irregular variables, like the distinction between sudden, continuous and meandering period changes, are tools that prove useful in organizing the discussion, but they fail to provide a clear picture of the physics at stake. An abundant literature is dedicated to what causes changes of pulsation regime, most of the time blaming the occurrence of a thermal pulse, but little is known about how such changes actually develop. They clearly have to be seen in the context of non-linear dynamics, with light curves prone to display features such as humps or double-peaking and to become chaotic, but we are unable to answer simple questions such as: will R Dor, a nearby AGB star that has been extensively observed and studied (T. R. Bedding et al., 1998), see its light curve evolve from its current irregular regime to a more stable one? and if yes, when will that happen?
@@
In summary, the present study has significantly clarified the picture that had been sketched in Paper I. In particular it has offered a detailed description of the relation between stellar parameters and the properties of the light curves of the Table 1 (Mno) sample and it has revealed a clear distinction between the curves of the Table 3 and Table 4 samples, which had been previously unnoticed. It has suggested that their evolution follows two different paths in the parameter space, the former possibly implying a transition out of the regime of stable pulsation. However, it has raised many questions which remain unanswered and the interpretations that it has suggested are far from being ascertained. It should serve as a guideline for future observations and analyses, which would hopefully help with progressing with the general understanding of the issues that have been addressed. It makes one feel that light curves still carry messages and information that we have not yet been able to decrypt.
@@ acknowledgments
To the extent that the analyses presented in the present article may have contributed some progress, the credit belongs to the innumerable observers around the world and to the AAVSO, without whom none of the results could have been obtained. We are accordingly deeply indebted to them: to the many observers for the high quality of their observations and to the AAVSO for the outstanding handling and reduction of the data that makes their use particularly easy and efficient. We are deeply grateful to Pr Tim Bedding for his interest in our work and for useful comments. Financial support from the World Laboratory, the Rencontres du Vietnam and the Vietnam National Space Center is gratefully acknowledged. This research is funded by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 103.99-2024.36.

### v2

@@ 3.5 Humpy curves and hump-less curves
In Paper I, we remarked that when a hump was observed on the ascending branch, there was a relation between its location and the parameters , and : namely the closer it was to minimal light, the larger were the values of these three parameters. Figure 8 illustrates the present situation. We selected curves displaying a clear hump on the ascending branch and classified them as type b, c or d according to the hump location: b for humps close to minimal light, d for humps close to maximal light and c in between. In the definitions of both what we call a strong hump and of which class the curve belongs to, there is of course some subjectivity which must be kept in mind. Moreover, what we define as humps are rather abrupt changes of the rate of increase of the luminosity: while c types are real humps, b and d types are rather flattening of the light minimum and maximum, respectively. Figure 8 reveals a number of remarkable features: all d types are from the Table 4 sample and, with the exception of V Cyg, all b types are from the Table 3 sample; instead, c types are present in all three samples and span a very large range of values; the general trends observed in Paper I are clearly confirmed and complemented with the dependence of the hump type on , and . Such humps may simply be the trivial effect of non-linear dynamics, similar humps are observed in RR Lyrae stars and in bump Cepheids with the Hertzsprung progression. In such a case, they may not be amenable to a simple interpretation, such as tracking a shock wave as suggested by L. S. Kudashkina & G. M. Rudnitskij (1994). Yet, their distribution in the parameter space suggests that they carry some additional information, which we however fail to decrypt.
@@ 4.1 Hot Bottom Burners
R Cen and R Nor have been identified as clear Hot Bottom Burning (HBB) candidates at the beginning of the TP-AGB by D. A. García-Hernández et al. (2013) and S. Uttenthaler et al. (2011, 2012) on the basis of detailed evaluations of the abundances of technetium, lithium, rubidium and zirconium in the spectra: they are Tc-poor and Li-rich and have low mass-loss rates at the level of 10-7 M⊙ yr-1, properties which they share with two other clear HBB candidates, SV Cas and RU Cyg, both semi-regular variables. They are expected to have initial masses of 4 to 6 solar masses, in contrast to more massive HBB candidates with very high mass-loss rates (10-5 M⊙ yr-1), known as OH/IR stars, at the end of the TP-AGB (D. A. García-Hernández et al., 2006, 2007, 2009). Their light curves are double-peaked (Figure 9) and that of R Cen has recently experienced a sudden change in period and amplitude. Moreover, they are spectacular outliers in colour vs period diagrams (S. Uttenthaler et al., 2019; P. Merchán Benítez et al., 2023; P. Whitelock et al., 2000). They have been commented upon in numerous earlier studies, both for R Cen (P. C. Keenan & J. Landi Dessy, 1966; G. Hawkins et al., 2001; W. S. G. Walker & J. Greaves, 2001; M. R. Templeton et al., 2005; J. R. Garcia & S. A. Dominguez, 2006; N. Vogt et al., 2016; B. Rea & J. C. Martin, 2024) and for R Nor (N. Vogt et al., 2016). Most authors take it as granted that the sudden changes observed in the light curve of R Cen are the result of a recent thermal pulse, as had been first proposed by P. R. Wood & D. M. Zarro (1981). Measurements of the radial velocity in the near infrared (T. Lebzelter et al., 2005) have given evidence for a clear anti-correlation with the light curve.
@@ 4.2 Mno outliers
The curve of W Dra is only marginally different from those of the Table 1 sample. While it was reasonable to exclude it from the studied sample, it is now clear that it might as well have been included, which would have had essentially no impact on the conclusions of the analysis. We did not find in the published literature evidence for the star to display abnormal features.
@@ 5 Discussion and conclusion
The emphasis in the present article has been to try to reveal general relations between the evolution of stars along the AGB and the parameters of their light curves, at the price of accepting some loss of rigour. For example, by using mean parameters , and , we lost the information contained in the differences between the members of the relevant pairs. Also, we accepted to ignore the differences, often significant, between the uncertainties attached to the parameters that have been used. This must be kept in mind when considering the results obtained and possibly extending the present work to future observations or analyses. By using five parameters to characterize the main features of a light curve, we have been able to clarify the picture that had been sketched in Paper I. Figure 10 illustrates the main results, which we summarize below. Table 7 lists, for each of the Table 1, 3, 4 and 5 samples, the mean values and rms deviations with respect to the mean of the curve parameters. Table 8 lists, for cases where a significant correlation has been observed between and another parameter, the result of a linear fit and the rms deviation of with respect to it. In three cases ( vs , and ) the linear fit is made jointly to Table 3 and Table 4 curves. The linear fits are illustrated in Figure 11.
When on the early part of the AGB, stars expand, cool down and tend to pulsate with increasing periods; when their period exceeds 100 days, they are classified as Mira variables and enter the sample considered in the present work. At that time, they are oxygen-rich with an M spectral type and have not yet experienced sufficient TDU events for their impact to be detectable. The sample of Mno curves listed in Table 1 corresponds to such stars. As the temperature of the star decreases and as its size increases, the light curve evolves from what was defined as first Mno family to what was defined as second Mno family in Paper I. Namely (black arrow in Figure 10) the period increases, the width parameter decreases, the phase at minimal light increases, weak humps may appear on the ascending branch close to minimal light, the pulsations become less regular, the colour indices increase and the ratio between the oscillation amplitudes in the infrared and in the visible increases.
@@
| Parameter | Sample | a | b | |
| vs 10 | Table 1 | 0.550 | 0.443 | 0.023 |
| vs 10 | Table 3 | 0.300 | 0.355 | 0.027 |
| vs | Tables 3+4 | 0.040 | 0.485 | 0.057 |
| vs | Tables 3+4 | 0.053 | 0.377 | 0.050 |
| vs | Tables 3+4 | 0.242 | 0.388 | 0.046 |
| vs | Table 5 | 0.111 | 0.743 | 0.047 |
| vs | Table 5 | 0.081 | 0.603 | 0.059 |
| vs | Table 5 | 0.320 | 0.557 | 0.064 |
@@
What the present work has made clear is that the transition from Mno to Myes, namely the occurrence of TDU events strong enough to have a detectable impact, occurs at very different stages along the TP-AGB. Indeed it is expected to never occur for stars of low initial masses (F. Herwig, 2005), and the curves of such stars stay in the Table 1 sample for the whole AGB life of the star. The presence in the Table 3 (0.3) sample of stars of Myes, and even S spectral types covering a broad spectrum of values makes a reliable interpretation in terms of their evolution along the AGB difficult. It suggests that TDU events may occur at very different stages on the TP-AGB; if such is the case, it would probably mean that they occur the earlier the larger the initial mass of the star, as shown by model simulations (F. Herwig, 2005). The curves of the Table 3 sample (blue arrow in Figure 10) seem to keep evolving in the same direction after the transition to the Myes and S spectral types: the periods and amplitudes of the oscillations keep increasing and they become less regular. This is consistent with the expectation that thermal pulses modify only progressively the internal state of the star. The absence of carbon-rich curves in the parameter space of the curves close to the more evolved curves of the Table 3 sample suggests that some of these curves leave the domain of stable pulsation and become chaotic before the transition from oxygen-rich to carbon-rich.
@@
Another major result obtained by the present work is the existence of curves leaving the Table 1 sample very early and transiting to the Myes and S spectral types with the width parameter increasing rather than decreasing. These are the curves of the Table 4 (0.3) sample and their location in the parameter space suggests that they will evolve to the Table 5 sample when the stars will become carbon-rich (red arrows in Figure 10). Here again, such a suggestion, albeit sensible and plausible, is far from being proven. The dispersion of the curve parameters about their mean is large for both the Table 4 and Table 5 samples, showing the complexity of the situation. In particular, the presence in the Table 5 sample of curves evolving with decreasing (possible successors of the Table 3 sample) and of curves evolving with increasing (possible successors of the Table 4 sample) contribute to such complexity. It is natural to speculate that the stars of the Table 4 sample have larger initial masses, and probably are closer to exhausting their convective layers than the stars of the Table 3 sample: the behaviour of the curves of HBB stars discussed in Subsection 4.1 gives support to such speculation as do the curves of stars that are known to be close to post-AGB, such as V Hya and VX Sgr.
@@
A major difficulty that has been recurrently met during the study is the lack of a clear picture of the physics governing the transition from a stable to an unstable regime of pulsation, namely of the nature of the stability strip in the Hertzsprung-Russell diagram. The distinction between Mira, semi-regular and irregular variables, like the distinction between sudden, continuous and meandering period changes, are tools that prove useful in organizing the discussion, but they fail to provide a clear picture of the physics at stake. An abundant literature is dedicated to what causes changes of pulsation regime, most of the time blaming the occurrence of a thermal pulse, but little is known about how such changes actually develop. They clearly have to be seen in the context of non-linear dynamics, with light curves prone to display features such as humps or double-peaking and to become chaotic, but we are unable to answer simple questions such as: will R Dor, a nearby AGB star that has been extensively observed and studied, see its light curve evolve from its current irregular regime (T. R. Bedding et al., 1998) to a more stable one? and if yes, when will that happen?
@@
In summary, the present study has significantly clarified the picture that had been sketched in Paper I. In particular it has offered a detailed description of the relation between stellar parameters and the properties of the light curves of the Table 1 (Mno) sample and it has revealed a clear distinction between the curves of the Table 3 and Table 4 samples, which had been previously unnoticed. It has suggested that their evolution follows two different paths in the parameter space, the former possibly implying a transition out of the regime of stable pulsation. However, it has raised many questions which remain unanswered and the interpretations that it has suggested are far from being ascertained. It should serve as a guideline for future observations and analyses, which would hopefully help with progressing with the general understanding of the issues that have been addressed. In particular, a detailed study of the shape of the ascending branches of the light curves would probably help with clarifying some of the issues that are still unclear and the consideration of light curves from different sources, such as OGLE and ASAS-SN, would be welcome. It makes one feel that light curves still carry messages and information that we have not yet been able to decrypt.
@@ acknowledgments
To the extent that the analyses presented in the present article may have contributed some progress, the credit belongs to the innumerable observers around the world and to the AAVSO, without whom none of the results could have been obtained. We are accordingly deeply indebted to them: to the many observers for the high quality of their observations and to the AAVSO for the outstanding handling and reduction of the data that makes their use particularly easy and efficient. We very much thank the anonymous referee for pertinent comments that helped with improving the presentation of the article. We are deeply grateful to Pr Tim Bedding for his interest in our work and for useful comments. Financial support from the World Laboratory, the Rencontres du Vietnam and the Vietnam National Space Center is gratefully acknowledged. This research is funded by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 103.99-2024.36.